One side of a rectangular stage is 2 meters longer than the other. If the diagonal is 10 meters, then what are the lengths of the sides?
Found 2 solutions by edjones, solver91311:
Answer by edjones(7569)

Let length of one side be x and the other x+2.
a^2+b^2=c^2
x^2+(x+2)^2=10^2
x^2+x^2+4x+4=100
2x^2+4x-96=0
x^2+2x-48=0
(x+8)(x-6)=0
x=-8, x=6
The sides are 8 and 6.
.
Ed

Answer by solver91311(16877)





Solve for


John